find the solution to the interpolation problem of finding a polynomial q(x) withdegree 2  and such that

q(x0)=y0

q(x1 )=y1

q'(x_1) = y'_1 ​ with x_0 < x_1x 0 ​

Under what exact conditions is \deg(q) = 2deg(q)=2?
Answer #1

Please comment if you need any clarification.

If you find my answer useful please put thumbs up. Thank you.

let q6x)= at batch Using given condition 9680)=40 967.) 4, 9.6x1) = 4, = at bilgt cao atbx; + C7, 2 Yo = 9Exo) Oude Y = 9(2)We have a- .1 Osting (4,-4,*,)(40-20° + x, C4-440) (no-2 b: y (xo-,) - 2 (40-46%) xi (26-2) (Yo-Yino) of no = 9; = 0 (al y,